a body travelling with uniform acceleration crosses two points A and B with velocities 30m/s and 40m/s respectively.Find speed of the body at mid point of A and B

Grade:11

1 Answers

Saurabh Kumar
8 years ago
Use, V2 = u2 + 2aS
Here, V = 40 m/s and u = 30 m/s, a is uniform and S is the distance between point A and B.
Let V’ is the velocity of the body at mid point, i.e. S/2
Then use V’2 = u2+ 2a (S/2)